OpenStudy (kaylala):

Find the Exact Value of: sin 5π/6 by using π/2 and π/3

OpenStudy (kaylala):

@mustafa2014

OpenStudy (anonymous):

\[\sin(\frac{ 5\pi }{ 6 })=\sin(\frac{ \pi }{ 3 }+\frac{ \pi }{ 2 })\]

OpenStudy (anonymous):